# Better Alignment with Instruction Back-and-Forth Translation

### v1

@@ Abstract
We propose a new method, instruction back-and-forth translation, to construct high-quality synthetic data grounded in world knowledge for aligning large language models (LLMs). Given documents from a web corpus, we generate and curate synthetic instructions using the backtranslation approach proposed by Li et al. (2023a), and rewrite the responses to improve their quality further based on the initial documents. Fine-tuning with the resulting (backtranslated instruction, rewritten response) pairs yields higher win rates on AlpacaEval than using other common instruction datasets such as Humpback, ShareGPT, Open Orca, Alpaca-GPT4 and Self-instruct. We also demonstrate that rewriting the responses with an LLM outperforms direct distillation, and the two generated text distributions exhibit significant distinction in embedding space. Further analysis shows that our backtranslated instructions are of higher quality than other sources of synthetic instructions, while our responses are more diverse and complex than those obtained from distillation. Overall we find that instruction back-and-forth translation combines the best of both worlds—making use of the information diversity and quantity found on the web, while ensuring the quality of the responses which is necessary for effective alignment.
@@ 1 Introduction
Given the same data quantity, fine-tuning Llama-2-70B on the instruction-response pairs from our data generation pipeline improves the AlpacaEval win rate by 3.6% compared to using the backtranslation data from previous work Li et al. (2023a), and by at least 3.2% compared to using other existing distillation datasets such as OpenOrca Lian et al. (2023), ShareGPT Chiang et al. (2023), Alpaca-GPT4 Peng et al. (2023) and Self-instruct Wang et al. (2022a) (Section 4). By asking a model to rewrite responses based on initial texts extracted from Dolma, we obtain a distribution of responses that interpolates between the original web text distribution and the distribution of outputs distilled from the same model. Fine-tuning on the rewritten responses in turn outperforms fine-tuning on the distilled responses corresponding to the same instructions (Section 5.1). In addition, we offer some insights into how instruction backtranslation and response rewriting affect the quality of instructions and responses respectively, especially in comparison with existing data generation methods (Sections 5.2 and 5.3). Overall our results suggest that back-and-forth translation offers an effective way to generate instruction-tuning data enriched with diverse information found on the web, while ensuring the quality of the response annotations by having aligned LLMs in the loop.
@@ 4 Fine-tuning results
Data source Data size Win rate (%) Dolma + filtering 12.8K 71.70 Dolma + rewriting 12.8K 73.44 Dolma + filtering + rewriting 12.8K 74.38 ClueWeb + filtering 12.8K 70.77 Open Orca 12.8K 74.20 ShareGPT 12.8K 72.69 Alpaca-GPT4 12.8K 71.17 Self-instruct 12.8K 65.11
@@
Data source Data size Win rate (%) Dolma + filtering 51.2K 87.42 Dolma + rewriting 51.2K 90.52 Dolma + filtering + rewriting 51.2K 91.74 ClueWeb + filtering 41.8K 88.18 Open Orca 51.2K 87.31 ShareGPT 51.2K 88.56 Alpaca-GPT4 51.2K 86.18 Self-instruct 51.2K 78.48 Dolma + filtering + rewriting 25.6K 90.22 Dolma + filtering + distilling 25.6K 87.58
@@ 6 Related Work
Most related to our approach is the work by Chen et al. (2023b), who train an LLM to generate both instructions and responses from web-scraped documents. In contrast to their method, we (i) generate instructions separately with backtranslation and then ask an LLM to improve the existing responses, (ii) obtain better performance with much fewer data (Table 5) (iii) generate more data (51.2K compared to 12.4K), (iv) offer more insights into the quality of our instructions and responses in comparison to other existing datasets.
@@
Some prior work studies characteristics of high-quality instruction-tuning data Liu et al. (2023) and proposes curation techniques accordingly. LIMA Zhou et al. (2024) carefully collects 1K fine-tuning samples via both internet sourcing and human annotation, and shows that strong performance can be achieved despite the small data quantity. Similarly, Chen et al. (2023a) demonstrates that performance gain is possible by fine-tuning on only a small subset of the original dataset (Alpaca), using ChatGPT as the quality evaluator. Zhao et al. (2024) finds that selecting only the 1K longest responses from existing datasets offers a very strong baseline, independent of GPT-4’s preference for longer texts. Evol-Instruct Xu et al. (2023) and Orca Mukherjee et al. (2023) manually prompt models to enhance the complexity of instructions, and subsequently, data generation (e.g. by asking for justification).
@@ Appendix D Rewriting ablations
As noted by the Alpaca team, length-controlled (LC) win rates alleviate length biases of GPT-4, but may favor models fine-tuned on its outputs. Our baselines of choice (described in Section 3.2) mostly distill outputs from GPT4 (in the case of Alpaca-GPT4 and Open Orca) and GPT-related models (ChatGPT for ShareGPT and GPT-3 for Self-instruct).
@@
Dolma + filtering 51.2K 16.71 11.90 17.66 9.63 Dolma + rewriting 51.2K 17.44 12.66 21.54 15.34 Dolma + filtering + rewriting 51.2K 16.68 12.07 19.09 12.86 ClueWeb + filtering 41.8K 15.93 11.89 18.39 10.26 Open Orca 51.2K 16.96 11.82 16.35 9.63 ShareGPT 51.2K 18.42 11.62 19.14 9.81 Alpaca-GPT4 51.2K 17.19 11.48 17.04 8.63 Self-instruct 51.2K 12.20 7.09 13.24 6.58 Dolma + filtering + rewriting 25.6K 17.90 13.05 17.55 11.86 Dolma + filtering + distilling 25.6K 17.80 11.95 17.50 10.50
@@ Appendix F Other NLP evaluations
Data source Data size HellaSwag ARC PIQA MMLU 7B scale Dolma + filtering 12.8K 60.6 59.1 76.1 35.6 Dolma + rewriting 12.8K 56.4 56.8 74.6 33.8 Dolma + filtering + rewriting 12.8K 56.8 56.7 74.5 36.6 ClueWeb + filtering 12.8K 60.3 57.5 75.7 35.1 Open Orca 12.8K 57.3 59.1 76.0 45.7 ShareGPT 12.8K 56.3 59.2 75.1 40.4 Alpaca-GPT4 12.8K 57.3 57.7 74.7 35.0 Self-instruct 12.8K 56.7 55.7 74.9 35.2 70B scale Dolma + filtering 51.2K 68.1 69.4 81.8 62.8 Dolma + rewriting 51.2K 66.4 69.3 81.2 61.9 Dolma + filtering + rewriting 51.2K 66.8 69.1 81.4 61.4 ClueWeb + filtering 41.8K 68.8 68.4 81.6 63.3 Open Orca 51.2K 65.8 66.6 78.6 61.8 ShareGPT 51.2K 65.7 68.4 79.7 60.0 Dolma + filtering + rewriting 25.6K 66.8 69.7 82.2 62.7 Dolma + filtering + distilling 25.6K 65.7 70.9 82.0 60.8

### v2

@@ Abstract
We propose a new method, instruction back-and-forth translation, to construct high-quality synthetic data grounded in world knowledge for aligning large language models (LLMs). Given documents from a web corpus, we generate and curate synthetic instructions using the backtranslation approach proposed by Li et al. (2023a), and rewrite the responses to improve their quality further based on the initial documents. Fine-tuning with the resulting (backtranslated instruction, rewritten response) pairs yields higher win rates on AlpacaEval than using other common instruction datasets such as Humpback, ShareGPT, Evol-Instruct, Open Orca, Alpaca-GPT4 and Self-instruct. We also demonstrate that rewriting the responses with an LLM outperforms direct distillation, and the two generated text distributions exhibit significant distinction in embedding space. Further analysis shows that our backtranslated instructions are of higher quality than other sources of synthetic instructions, while our responses are more diverse and complex than those obtained from distillation. Overall we find that instruction back-and-forth translation combines the best of both worlds—making use of the information diversity and quantity found on the web, while ensuring the quality of the responses which is necessary for effective alignment.
@@ 1 Introduction
Given the same data quantity, fine-tuning Llama-2-70B on the instruction-response pairs from our data generation pipeline improves the AlpacaEval win rate by 3.6% compared to using the backtranslation data from previous work Li et al. (2023a), and by at least 3.2% compared to using other existing distillation datasets such as OpenOrca Lian et al. (2023), ShareGPT Chiang et al. (2023), Evol-Instruct Xu et al. (2023), Alpaca-GPT4 Peng et al. (2023) and Self-instruct Wang et al. (2022a) (Section 4). By asking a model to rewrite responses based on initial texts extracted from Dolma, we obtain a distribution of responses that interpolates between the original web text distribution and the distribution of outputs distilled from the same model. Fine-tuning on the rewritten responses in turn outperforms fine-tuning on the distilled responses corresponding to the same instructions (Section 5.1). In addition, we offer some insights into how instruction backtranslation and response rewriting affect the quality of instructions and responses respectively, especially in comparison with existing data generation methods (Sections 5.2 and 5.3). Overall our results suggest that back-and-forth translation offers an effective way to generate instruction-tuning data enriched with diverse information found on the web, while ensuring the quality of the response annotations by having aligned LLMs in the loop.
@@ 3.2 Baselines
  • Evol-Instruct Xu et al. (2023): starting from the Alpaca instruction set, this work uses a set of evolution prompts to rewrite the instructions to improve their complexity. This evolution process is repeated multiple times, using an eliminator in between to filter out the failed instructions. Responses are then distilled from ChatGPT.
@@ 4 Fine-tuning results
Data source Data size Win rate (%) Dolma + filtering 12.8K 71.70 Dolma + rewriting 12.8K 73.44 Dolma + filtering + rewriting 12.8K 74.38 ClueWeb + filtering 12.8K 70.77 Open Orca 12.8K 74.20 ShareGPT 12.8K 72.69 Evol-Instruct 12.8K 72.32 Alpaca-GPT4 12.8K 71.17 Self-instruct 12.8K 65.11
@@
Data source Data size Win rate (%) Dolma + filtering 51.2K 87.42 Dolma + rewriting 51.2K 90.52 Dolma + filtering + rewriting 51.2K 91.74 ClueWeb + filtering 41.8K 88.18 Open Orca 51.2K 87.31 ShareGPT 51.2K 88.56 Evol-Instruct 51.2K 86.05 Alpaca-GPT4 51.2K 86.18 Self-instruct 51.2K 78.48 Dolma + filtering + rewriting 25.6K 90.22 Dolma + filtering + distilling 25.6K 87.58
@@ 6 Related Work
Most related to our approach is the work by Chen et al. (2023b), who train an LLM to generate both instructions and responses from web-scraped documents. In contrast to their method, we (i) generate instructions separately with backtranslation and then ask an LLM to improve the existing responses, (ii) obtain better performance with much fewer data (Table 5) (iii) generate more data (51.2K compared to 12.4K), (iv) offer more insights into the quality of our instructions and responses in comparison to other existing datasets. In addition, concurrent work by Zheng et al. (2024) also proposes more detailed scoring and refinement prompts to improve the instruction curation and response formatting of the backtranslation pipeline from Li et al. (2023a), applying it to Chinese text data.
@@
Some prior work studies characteristics of high-quality instruction-tuning data Liu et al. (2023) and proposes curation techniques accordingly. LIMA Zhou et al. (2024) carefully collects 1K fine-tuning samples via both internet sourcing and human annotation, and shows that strong performance can be achieved despite the small data quantity. Similarly, Chen et al. (2023a) demonstrates that performance gain is possible by fine-tuning on only a small subset of the original dataset (Alpaca), using ChatGPT as the quality evaluator. Zhao et al. (2024) finds that selecting only the 1K longest responses from existing datasets offers a very strong baseline, independent of GPT-4’s preference for longer texts. Evol-Instruct Xu et al. (2023) and Orca Mukherjee et al. (2023) manually prompt models to enhance the complexity of instructions, and subsequently, data generation (e.g. by asking for justification). Fan et al. (2024) reformats the responses of existing instruction data to augment them with relevant information and align them with pre-determined criteria set by humans.
@@ Appendix D Rewriting ablations
As noted by the Alpaca team, length-controlled (LC) win rates alleviate length biases of GPT-4, but may favor models fine-tuned on its outputs. Our baselines of choice (described in Section 3.2) mostly distill outputs from GPT4 (in the case of Alpaca-GPT4 and Open Orca) and GPT-related models (ChatGPT for ShareGPT and Evol-Instruct, GPT-3 for Self-instruct).
@@
Dolma + filtering 51.2K 16.71 11.90 17.66 9.63 Dolma + rewriting 51.2K 17.44 12.66 21.54 15.34 Dolma + filtering + rewriting 51.2K 16.68 12.07 19.09 12.86 ClueWeb + filtering 41.8K 15.93 11.89 18.39 10.26 Open Orca 51.2K 16.96 11.82 16.35 9.63 ShareGPT 51.2K 18.42 11.62 19.14 9.81 Evol-Instruct 51.2K 15.65 9.51 17.89 8.94 Alpaca-GPT4 51.2K 17.19 11.48 17.04 8.63 Self-instruct 51.2K 12.20 7.09 13.24 6.58 Dolma + filtering + rewriting 25.6K 17.90 13.05 17.55 11.86 Dolma + filtering + distilling 25.6K 17.80 11.95 17.50 10.50
@@ Appendix F Other NLP evaluations
Data source Data size HellaSwag ARC PIQA MMLU 7B scale Dolma + filtering 12.8K 60.6 59.1 76.1 35.6 Dolma + rewriting 12.8K 56.4 56.8 74.6 33.8 Dolma + filtering + rewriting 12.8K 56.8 56.7 74.5 36.6 ClueWeb + filtering 12.8K 60.3 57.5 75.7 35.1 Open Orca 12.8K 57.3 59.1 76.0 45.7 ShareGPT 12.8K 56.3 59.2 75.1 40.4 Evol-Instruct 12.8K 54.2 56.1 75.3 35.9 Alpaca-GPT4 12.8K 57.3 57.7 74.7 35.0 Self-instruct 12.8K 56.7 55.7 74.9 35.2 70B scale Dolma + filtering 51.2K 68.1 69.4 81.8 62.8 Dolma + rewriting 51.2K 66.4 69.3 81.2 61.9 Dolma + filtering + rewriting 51.2K 66.8 69.1 81.4 61.4 ClueWeb + filtering 41.8K 68.8 68.4 81.6 63.3 Open Orca 51.2K 65.8 66.6 78.6 61.8 ShareGPT 51.2K 65.7 68.4 79.7 60.0 Evol-Instruct 51.2K 66.8 68.5 80.4 60.9 Dolma + filtering + rewriting 25.6K 66.8 69.7 82.2 62.7 Dolma + filtering + distilling 25.6K 65.7 70.9 82.0 60.8